OpenStudy (anonymous):

Find the linear approximation to f(x) at x = x o. Graph the function and its linear approximation. f(x) = sin 3x, x 0, = 0

OpenStudy (anonymous):

@electrokid can u help?

OpenStudy (anonymous):

@paddyfitz can u help?

OpenStudy (anonymous):

so, a linear approximation at \(x=x_0\) means, that if \(x\) changes by \(\Delta x\), estimate linearly, what \(f(x+\Delta x)\) would be.

OpenStudy (anonymous):

ok so i tried can u check it?

OpenStudy (anonymous):

nvm i think i did it wrong :/

OpenStudy (anonymous):

no, if x changes by dx, what is the change-> dy? \[f'(x)=3\cos(3x)\\\implies \Delta f(x_0)=3\cos(3x_0)\times\Delta x\\ f(x_0+\Delta x)=f(x_0)+\Delta f(x_0)\\ \qquad=\sin(3x_0)+3\cos(3x)\Delta x\]

OpenStudy (anonymous):

ohh ok what does the triangle represent?

OpenStudy (anonymous):

delta-> a small change in..

OpenStudy (anonymous):

ohh ok so

OpenStudy (anonymous):

following Leibnitz's notation \[\lim_{\Delta x\to0}\frac{\Delta y}{\Delta x}=\frac{dy}{dx}\]

OpenStudy (anonymous):

ohh okay so what would be the linear approx for this problem

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

infact, for any problem... try this out set \(f(x)=\ln(x)\), follow the same steps, to find \(\ln(0.1)\), use \(x_0=0, \Delta x=0.1\) and see how close you are

OpenStudy (anonymous):

or may be x=1.1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay i'm stuck

OpenStudy (anonymous):

haha. lets get you un-entangled. where you at?

OpenStudy (anonymous):

so will i start like this f(0) + f'(0) ( x - 0)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes. that is it

OpenStudy (anonymous):

but since \(\ln(0)\) is undefined, lets take "x=1" and "dx = 0.1"

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so can u help?